For the cell, Fe|Fe 2+ (0.01 M)|| Ag+ (0.1 M)|Ag, Write the cell reaction and calculate the emf of the cell at 298 K, if standard electrode potentials of Fe and Ag electrodes are -0.44 V and 0.8 V respectively. Calculate the change in standard free energy ∆Go for the reduction of 1 mole of Ag+.
1
Expert's answer
2019-09-10T03:47:26-0400

Solution.

"EMF = Ec - Ea"

Since the standard electrode potential of silver is greater than the potential of iron, silver will be the cathode, and iron the anode.

"C: Ag^+ + 1e = Ag^0"

"A: Fe - 2e = Fe^{2+}"

General reaction scheme:

"Ag^+ + Fe = Ag^0 + Fe^{2+}"

"Ec = Eo + \frac{0.059*lg(Ag^+)}{1}"

"Ea = Eo + \frac{0.059 \times lg(Fe^{2+})}{2}"

Ec = 0.741 V

Ea = -0.499 V

EMF = 1.24 V

EMFo = 1.24 V


"\Delta G = -z\times F\times Eo"

z = 1

"\Delta G = -119641,4 \frac{kJ}{mole}"

Answer:

"C: Ag^+ + 1e = Ag^0"

"A: Fe - 2e = Fe^{2+}"

"Ag^+ + Fe = Ag^0 + Fe^{2+}"

EMF = 1.24 V

"\Delta G = -119641,4 \frac{kJ}{mole}"